If marginal labour costs are given by the equation MLC = 20 + 8L. Calculate the cost of employing the first two labourers
Answer #1

MLC = 20 + 8L

When L = 2,

MLC = 20 + 8(2) = 36

The cost of employing the first two labourers = $36